The heat generated in a circuit is given by H = I2Rt joule where I is current, R is resistance and t is time. If the percentage errors in measuring I, R and t are 1%, 2% and 1% respectively, the maximum error in measuring heat will be

Grade:Upto college level

4 Answers

Saurabh Singh
10 years ago
( error in H) /H= 2*(error in I)/I +( error in R)/R +( error in t)/t

%error in H = 2*1+2+1
=5%

Thanks & Regards

Saurabh Singh,

askIITians Faculty

B.Tech.

IIT Kanpur

Kamlesh Kumar
10 years ago
MAXIMUM ERROR IN MEASURING HEAT

?H/H X 100 = 2 X ?I / I X 100 + ? R / R X 100 + ? T / T X 100

= 2 X 1 + 2 + 1 = 5

THANKS &REGARDS
kamlesh kumar
askiitians faculty

Rohit
5 years ago
Maximum error = error in Current+error in time +error in resistence
We know that Heat H=I^2RT
Then total error in heat =2×2+1+2=6% answer
Anushka
4 years ago
As we know heat=I^2RT
Therefore, the total error in heat% is=(2^2+1+1)
Ans=(4+1+1)=6%.
[NOTE:maximum error =error in current + error in resistance + error in time ]